OpenStudy (anonymous):

if you have a family of 4 children what is the probability of having 1 boy?

OpenStudy (sweetburger):

1/2 it does not matter how many children you have

OpenStudy (anonymous):

let me reword this.... you have a family of 4 children what is the probability that at least 1 is a boy?

OpenStudy (jonnyvonny):

Every 1 child, there was a chance to have either a boy or girl; so 1/2 per child, and that chance times itself 4 times, for you only have 1 boy out of 3.\[(1/2)*(1/2)*(1/2)*(1/2)= 1/8=.125=12.5\]

OpenStudy (jonnyvonny):

12.5%, wait, so you want ONLY 1 boy or at least 1?

OpenStudy (anonymous):

Is it one boy or at least one boy?..

OpenStudy (anonymous):

at least one boy

OpenStudy (anonymous):

P(at least 1 boy)=1-P(all girls)=1-1/2*1/2*1/2*1/2

OpenStudy (anonymous):

At least one boy: 15/16

OpenStudy (anonymous):

how did you get 15/16

OpenStudy (anonymous):

Look at eashy's answer

OpenStudy (jonnyvonny):

Are we including hermaphadites?...

OpenStudy (anonymous):

Not sure if you knew stef but this is actually a famous mathematical paradox.. Look up the Boy or Girl paradox

OpenStudy (anonymous):

eashy did you get 15/16 also?

OpenStudy (anonymous):

Simply put the chance is 15/16 because the only chance that the children wouldn't include a boy is if they were all girls.

OpenStudy (anonymous):

P(at least one boy)+P(all girls)=1

OpenStudy (anonymous):

Got it Thank You :)